OpenStudy (anonymous):

Simplify the following expression: 2x - 2y + 5z - 2x - y + 3z

OpenStudy (anonymous):

2x-2x=

OpenStudy (anonymous):

I think the equation would be... 2x - 2y + 5z - 2x - y + 3z = (2x - 2x) + (-2y - y) + (5z + 3z) = ?

OpenStudy (anonymous):

@shannayoung